OpenStudy (anonymous):

Solve each inequality or equation. Graph the solution (number line). 5 abs. value 3w+2 abs value -3 >7 (I don't know how to do the absolute value sign so I typed it down...

OpenStudy (anonymous):

\[5\left| 3w \right|+2\left| -3 \right|>7\]?

OpenStudy (anonymous):

5|3w+2|-3>7 (FOund the sign...)

OpenStudy (anonymous):

Ok. Well, you need to add three to both sides and then divide by five to isolate the absolute value.

OpenStudy (anonymous):

5|3w+2|>10 Do you divide the 5 to both sides or do you distribute it?

OpenStudy (anonymous):

Divide :)

OpenStudy (anonymous):

3/5w +2/5> 10 or 3/5w+2/5<-10 3/5w>9 and 3/5 3/5w<-10.4 w>16 and w<17 and 1/3 Is this correct?

OpenStudy (anonymous):

\[\left| 3w+2 \right|=2\]

OpenStudy (anonymous):

sorry, greater than

OpenStudy (anonymous):

2

OpenStudy (anonymous):

For this problem i need two answers so sorry but can you point blanky just tell me the answer and how you got there on the margin next to it, please. <^.^>

OpenStudy (anonymous):

3/5w +2/5> 2 or 3/5w+2/5<-2 3/5w>1 and 3/5 3/5w<-2 and 2/5 w>2/3 w<-4 Is this correct? (Sorry about the earlier comment I was just confused what you were telling me)

OpenStudy (anonymous):

w is greater than zero or less than negative four thirds. I can't do the number line thing, sorry. You have the general idea, but I think your execution is a little shoddy. The absolute value breaks down to \[3w+2>2\]or\[3w+2<-2\]You solve for w in both equations by first substracting 2 from both sides\[3w >0\]or\[3w<-4\]Then you need to divide both sides by three, which will give the answers I showed you. Always remember to test your values by tesing near values which you think don't make the inequality true

OpenStudy (anonymous):

Thank you so much! <^.^>